# Bayesian Characterization and Identifiability of Multilevel Quantum Systems

### v1

@@ 6.2 Adaptive Pulse Parameterization
500 4.0108 (9.260E-07) 127.8 (3.645E-06) 43.242 (2.592) 26.196 (1.719)
1000 4.0108 (6.813E-07) 127.8 (2.469E-06) 44.185 (2.304) 24.904 (1.928)
$\infty$ 4.0108 (8.324E-07) 127.8 (2.686E-06) 44.951 (2.416) 24.171 (1.494)
Figure 8 shows the distribution of the parameters and prediction errors whose mean and standard deviation are presents in Tab. 5. Firstly, we see that in the absence of shot noise, the parameter mean is better estimated for all four parameters, whereas the presence of shot noise results in a shift in mean value and a larger spread. However, the distribution (and the statistical moment) of the error shows that an adaptive pulse leads to an order of magnitude reduction in error mean and variance (in the absence of shot noise) when compared to the fixed pulse parameters; the mean errors are also an order of magnitude smaller with comparable variances on the order of $10^{-3}$. This suggests that the BExD approach is able to reliably estimate the model parameters, even though the shot noise significantly affects the accuracy of the estimated parameters Furthermore, despite severe undersampling (using only 500 shots), the errors in the predictions from the calibrated model are well within typical bounds $\approx 99.5\%$ [41] for several applications of NISQ testbeds.
Figure 8: (Top) Converged parameter distribution when performing noisy experiments and (Bottom) distribution of the error in the population when using different number of shots. Note the parameter distributions are translated by the mean of the exact distributions.
Table 5: Mean and standard deviation (in the brackets) of the error using an adaptive pulse parameterization
@@
500 3.85E-05 (2.29E-03) -3.25E-05 (1.57E-03) -6.09E-06 (2.18E-03)
1000 7.71E-06 (5.87E-04) 8.62E-07 (4.15E-04) -8.57E-06 (5.70E-04)
$\infty$ -1.76E-06 (2.76E-04) 4.87E-06 (1.93E-04) -3.11E-06 (2.57E-04)
@@ 7 Conclusion
The conventional techniques for quantum characterization, whether probabilistic or deterministic, rely on measurements from a set of experiments that are designed a-priori. These measurements are then used to perform an offline calibration of the mathematical model to reconstruct the measurements from the experimental testbed. In most cases, the theoretical identifiability of the model parameters is neglected. Even when the identifiability conditions are known, manual design of experiments satisfying such conditions is not trivial. Hence, this work presented an automatic, online Bayesian framework for characterizing multilevel open quantum systems described by the Lindblad master equation. The approach automatically identifies optimal robust experiments that provide maximum information on the model parameters to be estimated. Furthermore, we proved the theoretical identifiability of the Schroedinger and Lindblad equations given a measured observable, and derived the sufficient conditions for the quantum state to establish the global, unique identifiability of the model parameters.
We demonstrated our approach to identify the frequencies and decoherence parameters on a numerical test problem subjected to both parameter and shot noise. The characterization study demonstrated that different parameters exhibit different rates of convergence and are dependent on both the pulse parameterization and magnitude of shot noise. It was seen that accurate estimation of the frequencies required shortened pulse sequences whereas estimation of decoherence parameters required longer pulse sequences. We introduced an adaptive pulse parameterization to improve the convergence of all estimates. The results show that the Bayesian experiment design framework with adaptive pulse parameterization is able to accurately estimate all four parameters in only 500 epochs (i.e. 500 different experiments). The approach is better able to estimate the distribution of the system frequencies than the decoherence parameters. Desipite this, the BExD approach yielded robust estimates of the parameters such that the mean prediction error and variances were on the order of $10^{-4}$. The results showed that the online BExD approach is able to design robust, optimal experiments to accurately characterize the quantum system described by the Lindblad master equation. Ongoing and future work includes experimentally demonstrating the technique on multi-qubit systems.

### v2

@@ 6.2 Iterative Pulse Extension
500 4.0108 (9.260×$\times 10^{-7}$) 127.8 (3.645×$\times 10^{-6}$) 43.242 (2.592) 26.196 (1.719)
1000 4.0108 (6.813×$\times 10^{-7}$) 127.8 (2.469×$\times 10^{-6}$) 44.185 (2.304) 24.904 (1.928)
$\infty$ 4.0108 (8.324×$\times 10^{-7}$) 127.8 (2.686×$\times 10^{-6}$) 44.951 (2.416) 24.171 (1.494)
Figure 8 shows the distribution of the parameters and prediction errors whose mean and standard deviation are presents in Tab. 5. Firstly, we see that in the absence of shot noise, the parameter mean is better estimated for all four parameters, whereas the presence of shot noise results in a shift in mean value and a larger spread. However, the distribution (and the statistical moment) of the error shows that iterative pulse extensions lead to an order of magnitude reduction in error mean and variance (in the absence of shot noise) when compared to the fixed pulse parameters; the mean errors are also an order of magnitude smaller with comparable variances on the order of $10^{-3}$. This suggests that the BExD approach is able to reliably estimate the model parameters, even though the shot noise significantly affects the accuracy of the estimated parameters Furthermore, despite severe undersampling (using only 500 shots), the errors in the predictions from the calibrated model are well within typical bounds $\approx 99.5\%$ [46] for several applications of NISQ testbeds.
Figure 8: (Top) Converged parameter distribution when performing stochastic experiments and (Bottom) distribution of the error in the population when using different number of shots. Note the parameter distributions are translated by the mean of the exact distributions.
Table 5: Mean and standard deviation (in the brackets) of the error using iterative pulse extensions.
@@
500 3.85×$\times 10^{-5}$ (2.29×$\times 10^{-3}$) -3.25×$\times 10^{-5}$ (1.57×$\times 10^{-3}$) -6.09×$\times 10^{-6}$ (2.18×$\times 10^{-3}$)
1000 7.71×$\times 10^{-6}$ (5.87×$\times 10^{-4}$) 8.62×$\times 10^{-7}$ (4.15×$\times 10^{-4}$) -8.57×$\times 10^{-6}$ (5.70×$\times 10^{-4}$)
$\infty$ -1.76×$\times 10^{-6}$ (2.76×$\times 10^{-4}$) 4.87×$\times 10^{-6}$ (1.93×$\times 10^{-4}$) -3.11×$\times 10^{-6}$ (2.57×$\times 10^{-4}$)
@@ 7 Conclusion
The conventional techniques for quantum characterization, whether probabilistic or deterministic, rely on measurements from a set of experiments that are designed a-priori. These measurements are then used to perform an offline calibration of the mathematical model to reconstruct the measurements from the experimental testbed. In most cases, the theoretical identifiability of the model parameters is neglected. Even when the identifiability conditions are known, manual design of experiments satisfying such conditions is not trivial. Hence, this work presented an automatic, online Bayesian framework for characterizing open qutrit systems described by the Lindblad master equation. The approach automatically identifies optimal robust experiments that provide maximum information on the model parameters to be estimated. Furthermore, we proved the theoretical identifiability of the Schroedinger and Lindblad equations given a measured observable, and derived the sufficient conditions for the quantum state to establish the global, unique identifiability of the model parameters.
We demonstrated our approach to identify the frequencies and decoherence parameters on a numerical test problem subjected to both parameter uncertainty and shot noise. The characterization study demonstrated that different parameters exhibit different rates of convergence and are dependent on both the pulse parameterization and magnitude of shot noise. It was seen that accurate estimation of the frequencies required shortened pulse sequences whereas estimation of decoherence parameters required longer pulse sequences. We introduced an iterative pulse extension approach to improve the convergence of all estimates. The results show that the Bayesian experiment design framework with iterative pulse extensions is able to accurately estimate all four parameters in only 500 epochs (i.e. 500 different experiments). The approach is better able to estimate the distribution of the system frequencies than the decoherence parameters. Desipite this, the BExD approach yielded robust estimates of the parameters such that the mean prediction error and variances were on the order of $10^{-4}$. The results showed that the online BExD approach is able to design robust, optimal experiments to accurately characterize the quantum system described by the Lindblad master equation. Ongoing and future work includes experimentally demonstrating the technique on multi-qubit systems.
@@ Acknowledgements
Appendix A Robustness of Bayesian Experimental Design To Parameter Settings
Figure 9 shows the distribution of parameters and prediction errors from an ensemble (with different random initialization of 2000 particles) of six runs using the same control parameterization as those in Section 6.2. Here, the experiments are considered to be stochastic with parameter uncertainty but without shot noise. The distributions are similar to those in Fig. 8 and shows the estimation is robust to different initialization strategies.
Figure 9: (Top) Converged parameter distribution over an emsemble of six runs with when performing stochastic experiments with parameter uncertainty and (Bottom) distribution of the error in the population. Note the parameter distributions are translated by the mean of the exact distributions.
Figure 10 shows the distribution of parameters and prediction errors when using different number of Monte Carlo particles. Here, the experiments are again considered to be stochastic with parameter uncertainty but without shot noise. The distributions show that the frequency parameters are estimated more accurately and the decoherence times, with only a few number of samples. The under-resolved parameterization of the distribution (Eq. (24)) and estimation of expectations by Monte Carlo sampling (Eq. (23)), leads to this decrease in estimation accuracy, which in turn leads to an increase in prediction errors. However, this shows that the BExD approach does not require an overwhelming number of particle but a sufficient number to cover the parameter space.
Figure 10: (Top) Converged parameter distribution when performing stochastic experiments with parameter uncertainty and (Bottom) distribution of the error in the population when using different number of Monte Carlo particles. Note the parameter distributions are translated by the mean of the exact distributions.